# Thermalization rates and quantum Ruelle-Pollicott resonances: insights from operator hydrodynamics

### v1

@@ Abstract
In thermalizing many-body quantum systems without conservation laws such as ergodic Floquet and random unitary circuits, local expectation values are predicted to decay to their equilibrium values exponentially quickly. In this work we derive a relationship between said exponential decay rate $\overline{g}$ and the operator spreading properties of a local unitary evolution. A hydrodynamical picture for operator spreading allows us to argue that, for random unitary circuits, $\overline{g}$ is encoded by the leading eigenvalue of a dynamical map obtained by enriching unitary dynamics with dissipation, in the limit of weak dissipation. We argue that the size of the eigenvalue does not depend on the details of this weak dissipation (given mild assumptions on properties of the ergodic dynamics), so long as it only suppresses large operators significantly. Our calculations are based on analytical results for random unitary circuits, but we argue that similar results hold for ergodic Floquet systems. These conjectures are in accordance with existing results which numerically obtain quantum many-body analogues of classical Ruelle-Pollicott resonances [T. Prosen J. Phys. A: Math. Gen. 35 L737 (2002), T. Mori, arXiv:2311.10304] by studying unitary evolutions subject to weak dissipation.
@@
Question 2 was explored in classical dynamical systems, where $\bar{g}$ is related to Ruelle-Pollicott (RP) resonances Ruelle (1986); Pollicott (1985), which are given by poles of matrix elements of the analytically continued resolvent $R(z)=(z-U)^{-1}$, located at eigenvalues of $U$. Here, $U$ is the unitary Frobenius-Perron operator generating the discrete (Floquet) dynamics on the phase space densityGaspard (2005). These poles are difficult to compute exactly outside of simple models Hasegawa and Saphir (1992). A useful approach for computing these resonances numerically is to coarse-grain the phase space, so that $U$ becomes non-unitary Weber et al. (2000); Klus et al. (2015). Then $U$ can have eigenvalues $\{e^{-i\epsilon_{j}}\}$ within the unit circle. It has been shown in certain cases that the largest eigenvalue is “frozen” (independent of the coarse graining, in a certain limit), coinciding with the leading RP resonance which sets the decay rate of autocorrelations Nonnenmacher (2003); Faure and Roy (2006)
@@
More recently, various authors have studied quantum analogues of RP resonances, by applying quantum versions of “coarse-graining of phase space” Prosen (2002, 2004); Mori (2024); Znidaric (2024). Ref. Prosen, 2002, 2004 considered adding a projection $\mathbb{P}_{\ell}$ to an ergodic Floquet system, so that the Heisenberg dynamics of an operator $O$ is given by $\mathbb{P}_{\ell}\mathbb{U}_{F}$. Here $\mathbb{U}_{F}[\cdot]=U_{F}^{\dagger}[\cdot]U_{F}$ generates the Heisenberg dynamics under the Floquet unitary $U_{F}$. $\mathbb{P}_{\ell}$ is a dissipative superoperator that projects out operators with support greater than $\ell$. The resulting superoperator $\mathbb{P}_{\ell}\mathbb{U}_{F}$ has eigenvalues $\{e^{-i\epsilon_{j}}\}$ where $\{\epsilon_{j}\}$ are no longer all real. Numerically, the $\epsilon_{j}$ with smallest (negative) imaginary part, which governs the slowest decaying eigenvector, was shown to describe $\bar{g}$ obtained from the purely unitary dynamics generated by $U_{F}$. More recently, it was argued in Ref. Mori, 2024 that in ergodic Floquet systems, if one adds a small amount of dephasing dissipation (see Fig. 1) and takes the system size to infinity first before taking the dissipation rate to zero, then the resulting dissipative Floquet operator has eigenvalues $e^{-i\epsilon_{j}}$ with an $\epsilon_{j}$ with smallest negative real part that matches well with the decay of autocorrelation functions (this order of limits has also been studied in the context of “anomalous relaxation”Sá et al. (2022); García-García et al. (2023); Yoshimura et al. (2023)). Ref. Mori, 2024 also gave some physical intuition for why $\bar{g}$ can be encoded as an eigenvalue of the dissipative Floquet operator: the unitary dynamics spreads an operator ballistically until it grows to a steady-state size $x^{*}\sim\frac{1}{\gamma}$, where dissipation and unitary evolution balance out. Then the dissipation decreases the amplitude of this operator by $e^{-\gamma x^{*}}\sim e^{-\bar{g}}$ where $\bar{g}\sim\mathcal{O}(1)$.
@@
Hydrodynamical description of operator spreading.—We will first review the hydrodynamical description of operator spreading in random unitary circuits derived in Refs. von Keyserlingk et al., 2018; Nahum et al., 2018. We will assume that the circuit has a brickwork geometry built out of two-site gates, acting on sites with local Hilbert space dimension $q$. Any operator can be written as a sum of $\mathbb{Z}_{q}$ Pauli strings, generated by $\mathbb{Z}_{q}$ clock and shift operators (from this point onward we will simply refer to these as operator strings). Therefore, we only need to consider the evolution of operator strings under the random unitary dynamics.
An operator string $\sigma^{\mu}$ generically evolves into a sum of operator strings: $\sigma^{\mu}(t)=\sum_{\nu}c_{\nu}^{\mu}(t)\sigma^{\nu}$ where $\nu$ runs over operator strings and $c_{\nu}^{\mu}(t)$ are complex coefficients. One important insight from Refs. von Keyserlingk et al., 2018; Nahum et al., 2018 is that the evolution of an operator under a layer of random unitaries roughly depends only on the left and right endpoints of the operator (the sites farthes to the left and right that the operator acts non-trivially). Therefore, we can group together all operator strings with the same left and right endpoints. For simplicity, we will let the original operator be at the left end of the system, so we roughly only need to consider the movement of the right endpoints of operators. [^2] Choosing the operator to start at the center of the system and/or keeping track of correlations between the left and right endpoints will only modify the results in this paper by subleading $\frac{1}{\mathrm{poly}(t)}$ factors. We denote the density of operators with right endpoint at site $x$ by
@@
where we sum over operator strings $\sigma^{\nu}$ with right endpoint at $x$ [^3] We will use the same convention as in Ref. von Keyserlingk et al., 2018, where each unit cell (labeled by $x$) actually has two $q$ dimensional sites. It was shown in Ref. von Keyserlingk et al., 2018 that for random unitary circuits, $\overline{\rho_{R}^{\mu}}(x,t)$ takes the form of a propagating Gaussian with a peak moving at butterfly velocity $v_{B}$ that depends only on $q$ and variance growing linearly with $t$. This result was derived using the observation that $\overline{\rho_{R}}(x,t)$ satisfies the recurrence relation
@@
onto the vector indexed by integer $x\in[1,L]$ with elements $\overline{\rho_{R}}(x,t)$. $T(p)$ is a stochastic matrix: the elements in every column sum to 1. This is because the unitarity of the evolution guarantees that $\mathrm{Tr}(\sigma^{\mu}(t)\sigma^{\mu}(t)^{\dagger})$ is conserved, so $\sum_{s}\rho_{R}^{\mu}(s)=1$ for all $t$.
The fact that $T(p)$ is a stochastic matrix means that it has an eigenvalue 1, with multiplicity 1. For Floquet systems and modified random circuits, the stochastic $L\times L$ matrix $T$ describing the evolution of $\overline{\rho_{R}}(x,t)$ may be different. We only require that $T$ is quasi-diagonal and that $T_{x,y}=T_{x+1,y+1}$ for $\mathcal{O}(1)<x,y<\mathcal{O}(L)$. The latter uniformity assumption corresponds to the statement that once an operator has grown large, but is still significantly smaller than the system size, how it grows subsequently does not depend on its current size.
@@
We used the simple observation that the trace over any nontrivial operator string is zero. Note that $\overline{\rho}_{R}(n,t)\sim q^{2n}|c_{\mu}^{\mu}(t)|^{2}$ where $n$ is the extent of $\sigma^{\mu}$, so the autocorrelation function decays with the same rate as $\overline{\rho_{R}}(n,t)$. $\overline{\rho_{R}}(x,t)$ can be computed from $T(p)^{t}$ acting on a vector $|n\rangle$ whose only nonzero element is at site $n$. Therefore, to evaluate $\overline{\rho}_{n,t}$, we simply compute
$\overline{\rho}(n,t)=\langle n|T(p)^{t}|n\rangle=\sum_{k}\lambda_{k}^{t}\langle n|k_{R}\rangle\langle k_{L}|n\rangle$ (6)
@@
The above derivation of autocorrelation decay shows that there is no dependence on the observable $O$ or the initial state. Any local operator $O$ can be mapped onto another local operator by a finite number of layers of unitaries, so the long-time dynamics is unaffected by the choice of $O$. Furthermore, the computation of (1) in a product state is similar to the above computation; we simply need to also sum over $|c_{\mu}^{\nu}(t)|^{2}$ where $\sigma^{\nu}$ is a operator string for which the product state is an eigenstate. It turns out that this only changes the $\mathrm{poly}(t)$ factor sup ; the exponential decay rate remains the same.
More generally, as long as $T$ is stochastic, quasi-diagonal, and uniform, we expect that autocorrelation functions decay exponentially with a constant rate, up to times of order $L$. The decay rate given by subleading eigenvalues of $T$. In the above example, it is the second leading eigenvalue of $T$, but more generally there may be $\mathcal{O}(1)$ other eigenvalues (smaller than 1 but possibly larger than $e^{-\bar{g}}$) that also only appear only after $t\sim\mathcal{O}(L)$ because they are related to boundary condition of the biased walk at $x\sim L$. We give an example of this in the supplemental material sup .
We will now show that coarse-graining methods described earlier all have the effect of slightly modifying the matrix $T$. For the particular $T(p)$ in Eq. (4), they remove the eigenvalue 1 while, in appropriate limits, maintaining the eigenvalues around $4p(1-p)$. More generally, systems described by other choices of $T$, the coarse-graining methods remove the $\mathcal{O}(1)$ eigenvalues that appear at times $t\sim\mathcal{O}(L)$ so that the exponential decay observed for $t<\mathcal{O}(L)$ persists for all times. We will show that this implies that, for Floquet systems, the rate is encoded as an eigenvalue of the non-unitary generator of time translations.
@@
Although the superoperator $\mathbb{P}_{\ell}\mathbb{U}_{F}$ has non-orthonormal left and right eigenvectors, at some $t$, the decay rate will still be dominated by the eigenvalue with the largest modulus. In other words, at some value of $t$, the eigenvalue modulus overcomes any transient features determined by wavefunction overlaps [^5] For example, after times $t\sim\mathcal{O}(1)$, the largest eigenvalue $\lambda_{0}=1$ of $T(p)$ begins to dominate. It follows that $\mathbb{P}_{\ell}\mathbb{U}_{F}$ needs to have a dominant eigenvalue $\lambda$. Now, notice that $\lambda$ implicitly depends on $\ell$; however we do not expect it to depend strongly on $\ell$. The idea is that autocorrelators for $\mathbb{U}_{F}^{t}$ can be represented as a sum over histories of Pauli matrices. $(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}$ involves the same histories, except those involving operators larger than $\ell$ are discarded. However arguments like those in Rakovszky et al. (2022); von Keyserlingk et al. (2022) suggest the contributions discarded are subleading, and suppressed in $\ell$. This suggests that $\langle O\,(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}[O]\rangle$ and $\langle O\,\mathbb{U}_{F}^{t}[O]\rangle=\langle O(t)O\rangle$ decay at similar exponential rates, so that $\lim_{\ell\rightarrow\infty}\lambda=e^{-\overline{g}}$. This establishes that the leading eigenvalue of $\mathbb{P}_{\ell}\mathbb{U}_{F}$ approaches $e^{-\overline{g}}$ as $\ell\rightarrow\infty$, assuming we take the thermodynamic $L\rightarrow\infty$ limit first.
Another way to see the independence of $\ell$ and the convergence to $e^{-\bar{g}}$ is the following. We do not expect $\left|\langle O\,(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}[O]\rangle\right|$ to depend strongly on $\ell$ for sufficiently large $\ell$ because the dominant eigenvectors of the evolution with $P_{\ell}$ are operators near size $\ell$, and according to our uniformity assumption, operators at size $\mathcal{O}(1)<\ell<\mathcal{O}(L)$ all behave in the same way under the hydrodynamical description. We expect this decay rate to converge to $e^{-\bar{g}}$ because we can choose $\ell$ to be very large, i.e. $\ell=\frac{1}{2}L$. Then $\mathbb{P}_{\ell}$ does not affect the dynamics at all up to $t\sim\mathcal{O}(L)$. According to our arguments regarding how $P(\ell)$ modifies the eigenvalue structure of $T$, this means that the same rate $\bar{g}$ describes how $\left|\langle O\,(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}[O]\rangle\right|$ decays both at times up to $\mathcal{O}(L)$ (where $\mathbb{P}_{\ell}$ does nothing) and at times much larger than $\mathcal{O}(L)$ (from which we extract $\lambda=e^{-\bar{g}}$).
@@
Similar to the hard truncation, the soft truncation described by $P(\gamma)$ also removes the eigenvalue 1 of $T(p)$. Furthermore, in the above limits, $[P(\gamma)T(p)]_{x,y}\sim T(p)_{x,y}$ for $x,y\ll\frac{1}{\gamma}$. We can use our results for $P(\ell)T(p)$ to show that the eigenvalues of $P(\gamma)T(p)$ are upper bounded by $4p(1-p)$. Specifically, $P(\ell)T(p)$ has the same eigenvalues as $P(\ell)\tilde{T}(p)$ where $\tilde{T}(p)$ is a symmetric real matrix (see sup ). Next, we can define $\tilde{T}_{\mathrm{pr}}(p)$ which has a maximum eigenvalue close to $\sim 4p(1-p)$ by $\tilde{T}(p)=\tilde{T}_{\mathrm{pr}}(p)-e^{-\gamma L}p^{2}|L\rangle\langle L|$. Since $P(\gamma),\tilde{T}(p)$, and $\tilde{T}_{\mathrm{pr}}(p)$ are all Hermitian and positive definite, and the spectrum of $P(\gamma)\tilde{T}(p)$ is the same as that of $P(\gamma)\tilde{T}_{\mathrm{pr}}(p)$ up to corrections $\sim\mathcal{O}(e^{-\gamma L})$, the spectrum of $P(\gamma)\tilde{T}(p)$ is bounded from above by the product of the largest eigenvalues of $P(\gamma)$ and $\tilde{T}_{\mathrm{pr}}(p)$, which is $e^{-\gamma}4p(1-p)$. We observe numerical evidence that the leading eigenvalue of $P(\gamma)T(p)$ indeed approaches $4p(1-p)$ as $L\to\infty,\gamma\to 0$ (see Fig. 2).
@@
Furthermore, it is straightforward to see that the eigenvector $\overline{\rho_{R,k}}(x)$ of $P(\gamma)T(p)$ corresponding to eigenvalue $\lambda_{k}$ is peaked at size $-\frac{\log(\lambda_{k})}{\gamma}$, as argued by Ref. Mori, 2024. The eigenstate satisfies
$\begin{split}\lambda_{k}\overline{\rho_{R,k}}(x)&=e^{-\gamma x}2p(1-p)\overline{\rho_{R,k}}(x)+e^{-\gamma(x-1)}\overline{\rho_{R,k}}(x-1)\\ &+e^{-\gamma(x+1)}\overline{\rho_{R,k}}(x+1)\end{split}$ (9)
@@
Suppose that $\overline{\rho_{R,k}}(x)$ is a smooth function of $x$ (we will show that this is true in an approximation of $P(\gamma)T(p)$ later in this section; there, $\overline{\rho_{R,k}}$ has a Gaussian-like shape). Then when $\overline{\rho_{R,k}}(x)$ peaks, it flattens out, so $\overline{\rho_{R,k}}(x^{*})\sim\overline{\rho_{R,k}}(x^{*}-1)\sim\overline{\rho_{R,k}}(x^{*}+1)$. If we in addition take $\gamma\to 0$, (9) gives $\lambda_{k}=e^{-\gamma x^{*}}$, so $x^{*}=-\frac{\log{\lambda_{k}}}{\gamma}$. Assuming that the leading eigenvalue of $P(\gamma)T(p)$ approaches $4p(1-p)$ in the limit $L\to\infty,\gamma\to 0$, we find that indeed operators get spread to a size $\sim-\frac{\log(4p(1-p))}{\gamma}$. Because the operator size distribution get frozen, and operators do not grow to size $\sim L$, our argument from the previous section applies to this case as well. Even for Floquet systems, where $T$ may be different from $T(p)$ in (4), we expect the same overall picture to hold: dissipation leads to operators growing ballistically to size $x_{*}=O(1/\gamma)$ and becoming stuck there subsequently, with their norm decaying $e^{-\gamma x_{*}t}$ ever after. It follows from a similar argument to that near Eq. (8) that the dissipative Floquet operator $\mathbb{P}_{\gamma}\mathbb{U}_{F}$ has a leading eigenvalue with modulus $e^{-\bar{g}}$.
@@
Discussion.— Using an operator hydrodynamical picture, which is exact for RUCs and conjectured to work for ergodic Floquet circuits, we considered how operators grow in the presence of two very different forms of dissipation (hard versus soft truncation). We argued that at late times, a generic initial local operator tends towards a quasi-steady state with a profile peaked at size $x_{*}\sim\ell,1/\gamma$ in the hard, soft cases respectively (the latter result agrees with a prediction of Mori (2024)). In the RUC case, regardless of the choice of dissipation, the corresponding eigenvalue of the dissipative hydrodynamical process tends exactly to $e^{-\overline{g}}$ in the limit of weak dissipation, which controls the decay of autocorrelations. In the Floquet setting, we used the same operator spreading model to argue (less rigorously) that the leading eigenvalue of the dissipative many-body time translation generator $\mathbb{P}_{\ell}\mathbb{U}_{F}$ (or $\mathbb{P}_{\gamma}\mathbb{U}_{F}$) similarly determines the decay rate of autocorrelators in the weak dissipation limit.
We argued more generally that our results should be independent of the details of dissipation, provided it only suppresses large operators. This prediction comes from from the fact that our upper bound of the spectrum of $P(\gamma)T(p)$ does not depend on the particular form of $P(\gamma)$, as long as it is diagonal, is close to 1 for small $x$, and zero as $L\to\infty$.

### v2

@@ Abstract
In thermalizing many-body quantum systems without conservation laws, such as ergodic Floquet and random unitary circuits, local expectation values are predicted to decay to their equilibrium values exponentially quickly. In this work we derive a relationship between said exponential decay rate $\overline{g}$ and the operator spreading properties of a local unitary evolution. A hydrodynamical picture for operator spreading allows us to argue that, for random unitary circuits, $\overline{g}$ is encoded by the leading eigenvalue of a dynamical map obtained by enriching unitary dynamics with dissipation, in the limit of weak dissipation. We argue that the size of the eigenvalue does not depend on the details of this weak dissipation (given mild assumptions on properties of the ergodic dynamics), so long as it only suppresses large operators significantly. Our calculations are based on analytical results for random unitary circuits, but we argue that similar results hold for ergodic Floquet systems. These conjectures are in accordance with existing results which numerically obtain quantum many-body analogues of classical Ruelle-Pollicott resonances [T. Prosen J. Phys. A: Math. Gen. 35 L737 (2002), T. Mori, arXiv:2311.10304] by studying unitary evolutions subject to weak dissipation.
@@
Question 2 was explored in classical dynamical systems, where $\bar{g}$ is related to Ruelle-Pollicott (RP) resonances Ruelle (1986); Pollicott (1985). In classical dynamical systems, a unitary Frobenius-Perron operator $U$ generates the discrete (Floquet) dynamics on the phase space density. It was shown that after coarse-graining the phase space, $U$ becomes non-unitary, and has eigenvalues $\{e^{-\mathrm{i}\epsilon_{j}}\}$ within the unit circle. In certain cases, the largest eigenvalue becomes “frozen” (independent of the coarse graining), and coincides with the leading RP resonance which sets the decay rate of autocorrelations.
@@
More recently, various authors have studied quantum analogues of RP resonances, by applying quantum versions of “coarse-graining of phase space” Prosen (2002, 2004); Mori (2024); Znidaric (2024). Ref. Prosen, 2002, 2004 considered adding a projection $\mathbb{P}_{\ell}$ to an ergodic Floquet system, so that the Heisenberg dynamics of an operator $O$ is given by $\mathbb{P}_{\ell}\mathbb{U}_{F}$. Here $\mathbb{U}_{F}[\cdot]=U_{F}^{\dagger}[\cdot]U_{F}$ generates the Heisenberg dynamics under the Floquet unitary $U_{F}$. $\mathbb{P}_{\ell}$ is a dissipative superoperator that projects out operators with support greater than $\ell$. The resulting superoperator $\mathbb{P}_{\ell}\mathbb{U}_{F}$ has eigenvalues $\{e^{-\mathrm{i}\epsilon_{j}}\}$ where $\{\epsilon_{j}\}$ are no longer all real. Numerically, the $\epsilon_{j}$ with smallest (negative) imaginary part, which governs the slowest decaying eigenvector, was shown to describe $\bar{g}$ obtained from the purely unitary dynamics generated by $U_{F}$. More recently, it was argued in Ref. Mori, 2024 that in ergodic Floquet systems, if one adds a small amount of dephasing dissipation (see Fig. 1) and takes the system size to infinity first before taking the dissipation rate to zero, then the resulting dissipative Floquet operator has eigenvalues $e^{-\mathrm{i}\epsilon_{j}}$ with an $\epsilon_{j}$ with smallest negative real part that matches well with the decay of autocorrelation functions (this order of limits has also been studied in the context of “anomalous relaxation”Sá et al. (2022); García-García et al. (2023); Yoshimura and Sá (2023)). Ref. Mori, 2024 also gave some physical intuition for why $\bar{g}$ can be encoded as an eigenvalue of the dissipative Floquet operator: the unitary dynamics spreads an operator ballistically until it grows to a steady-state size $x^{*}\sim\frac{1}{\gamma}$, where dissipation and unitary evolution balance out. Then the dissipation decreases the amplitude of this operator by $e^{-\gamma x^{*}}\sim e^{-\bar{g}}$ where $\bar{g}\sim\mathcal{O}(1)$.
@@
Hydrodynamical description of operator spreading.—We will first review the hydrodynamical description of operator spreading in RUCs derived in Refs. von Keyserlingk et al., 2018; Nahum et al., 2018. We will assume that the circuit has a brickwork geometry built out of two-site gates, acting on sites with local Hilbert space dimension $q$. Any operator can be written as a sum of operator strings $\sigma^{\nu}$, which are products of local Hermitian basis operators (for $q=2$, these are just Pauli strings)Nahum et al. (2022). Therefore, we only need to consider the evolution of operator strings.
An operator string $\sigma^{\mu}$ generically evolves into a sum of operator strings: $\sigma^{\mu}(t)=\sum_{\nu}c_{\nu}^{\mu}(t)\sigma^{\nu}$ where $\nu$ runs over operator strings and $c_{\nu}^{\mu}(t)$ are complex coefficients. One important insight from Refs. von Keyserlingk et al., 2018; Nahum et al., 2018 is that the evolution of an operator under a layer of random unitaries, averaged over such random unitaries, roughly depends only on the left and right endpoints of the operator (the sites farthest to the left and right that the operator acts non-trivially).  [^2] In the present work, following Refs. Nahum et al. (2018); von Keyserlingk et al. (2018), we will work with quantities averaged over an ensemble of circuits (with averaging denoted by the overline) with the conjecture that the average accurately captures the behavior of a typical circuit in the ensemble.
Therefore, we can group together all operator strings with the same left and right endpoints. For simplicity, we will let the original operator be at the left end of the system, so we only need to consider the movement of the right endpoints of operators. [^3] Choosing the operator to start at the center of the system and/or keeping track of correlations between the left and right endpoints will only modify the results in this paper by subleading $\frac{1}{\mathrm{poly}(t)}$ factors. We denote the total weight on operators with right endpoint at site $R$, averaged over circuits, by
@@
where we sum over operator strings $\sigma^{\nu}$ with right endpoint at $x$ [^4] We will use the same convention as in Ref. von Keyserlingk et al., 2018, where each unit cell (labeled by $x$) actually has two $q$ dimensional sites. Since a single layer of local unitaries can change any Pauli-string to any other Pauli-string with the same support, the late time evolution does not depend on the particular choice of the initial operator $\sigma^{\mu}$, and we will drop this label. It was shown in Ref. von Keyserlingk et al., 2018 that for RUCs, $\overline{\rho_{R}^{\mu}}(x,t)$ takes the form of a propagating Gaussian with a peak moving at butterfly velocity $v_{B}$ that depends only on $q$, where the peak width grows as tproportional-$\propto\sqrt{t}$. This result was derived using the observation that $\overline{\rho_{R}}(x,t)$ satisfies the recurrence relation
@@
onto the vector indexed by integer $x\in[1,L]$ with elements $\overline{\rho_{R}}(x,t)$. $T(p)$ is a stochastic matrix: the elements in every column sum to 1. This is because the unitarity of the evolution guarantees that $\mathrm{Tr}(\sigma^{\mu}(t)\sigma^{\mu}(t)^{\dagger})$ is conserved, so $\sum_{s}\overline{\rho_{R}}^{\mu}(s)=1$ for all $t$.
The fact that $T(p)$ is a stochastic matrix means that it has an eigenvalue 1 (in this case with multiplicity 1). We conjecture that for Floquet systems, the operator dynamics is described by a similar stochastic matrix; the key properties we demand are that $T$ is quasi-diagonal and that $T_{x,y}=T_{x+1,y+1}$ for $\mathcal{O}(1)<x,y<\mathcal{O}(L)$. The latter uniformity assumption corresponds to the statement that once an operator has grown large, but is still significantly smaller than the system size, how it grows subsequently does not depend on its current size.
@@
We used the simple observation that the trace over any nontrivial operator string is zero. Note that $\overline{\rho_{R}}(n,t)\sim q^{2n}|c_{\mu}^{\mu}(t)|^{2}$ where $n$ is the extent of $\sigma^{\mu}$, so the autocorrelation function decays with the same rate as $\overline{\rho_{R}}(n,t)$. $\overline{\rho_{R}}(x,t)$ can be computed from $T(p)^{t}$ acting on a vector $|n\rangle$ whose only nonzero element is at site $n$. Therefore, to evaluate $\overline{\rho_{R}}(n,t)$, we simply compute
$\overline{\rho_{R}}(n,t)=\langle n|T(p)^{t}|n\rangle=\sum_{k}\lambda_{k}^{t}\langle n|k_{R}\rangle\langle k_{L}|n\rangle$ (6)
@@
The above derivation of autocorrelation decay shows that there is no dependence on the observable $O$ or the initial state. The intuition is that operator growth dynamics, at long times, does not depend on the specific structure of the initial local operator $O$. Likewise, the computation of (1) in a product state is similar to the above computation; we simply need to also sum over $|c_{\mu}^{\nu}(t)|^{2}$ where $\sigma^{\nu}$ is a operator string for which the product state is an eigenstate. It turns out that this only changes the $\mathrm{poly}(t)$ factor sup ; the exponential decay rate remains the same.
More generally, as long as $T$ is stochastic, quasi-diagonal, and uniform, we expect autocorrelation functions to decay exponentially with a constant rate, up to times of order $L$, at which point the operator has spread across the system. The decay rate is given by the subleading eigenvalues of $T$. In the above example, it is the second leading eigenvalue of $T$ [^6] More generally there may be other eigenvalues (smaller than 1 but possibly larger than $e^{-\bar{g}}$) that, like eigenvalue $1$, only contribute once $t\sim\mathcal{O}(L)$ because they are related to boundary condition of the biased walk at $x\sim L$. See sup for an example.
We will now show that coarse-graining methods described earlier all have the effect of slightly modifying the matrix $T$. For the particular $T(p)$ in Eq. (4), they remove the eigenvalue 1 while, in appropriate limits, maintaining only the eigenvalues around and below? $4p(1-p)$. More generally, for systems described by other choices of $T$, the coarse-graining method results in the exponential decay observed for $t<\mathcal{O}(L)$ persisting for all times. We will show that this implies that, for Floquet systems, the rate $\overline{g}$ is encoded as an eigenvalue of the non-unitary generator of time translations.
@@
Although the superoperator $\mathbb{P}_{\ell}\mathbb{U}_{F}$ has non-orthonormal left and right eigenvectors, at some $t$, the decay rate will still be dominated by the eigenvalue with the largest modulus. In other words, at some value of $t$, the eigenvalue modulus overcomes any transient features determined by wavefunction overlaps [^7] For example, after times $t\sim\mathcal{O}(L)$, the largest eigenvalue $\lambda_{0}=1$ of $T(p)$ begins to dominate. It follows that $\mathbb{P}_{\ell}\mathbb{U}_{F}$ needs to have a dominant eigenvalue $\lambda$. Now, notice that $\lambda$ implicitly depends on $\ell$; however we do not expect it to depend strongly on $\ell$. The idea is that autocorrelators for $\mathbb{U}_{F}^{t}$ can be represented as a sum over histories of Pauli matrices. $(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}$ involves the same histories, except those involving operators larger than $\ell$ are discarded. However arguments like those in Rakovszky et al. (2022); von Keyserlingk et al. (2022) suggest the contributions discarded are subleading, and suppressed in $\ell$. This suggests that $\langle O\,(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}[O]\rangle$ and $\langle O\,\mathbb{U}_{F}^{t}[O]\rangle=\langle O(t)O\rangle$ decay at similar exponential rates, so that $\lim_{\ell\rightarrow\infty}\lambda=e^{-\overline{g}}$. This establishes that the leading eigenvalue of $\mathbb{P}_{\ell}\mathbb{U}_{F}$ approaches $e^{-\overline{g}}$ as $\ell\rightarrow\infty$, assuming we take the thermodynamic $L\rightarrow\infty$ limit first.
@@
Similar to the hard truncation, the soft truncation described by $P(\gamma)$ also removes the eigenvalue 1 of $T(p)$. Furthermore, in the above limits, $[P(\gamma)T(p)]_{x,y}\sim T(p)_{x,y}$ for $x,y\ll\frac{1}{\gamma}$. We can use our results for $P(\ell)T(p)$ to show that the eigenvalues of $P(\gamma)T(p)$ are upper bounded by $4p(1-p)$. Specifically, $T(p)$ has the same spectrum as the symmetric matrix $\tilde{T}(p)$sup . Next, we can define $\tilde{T}_{\mathrm{pr}}(p)$ which has a maximum eigenvalue close to $\sim 4p(1-p)$ (according to the spectrum of $P(\ell)T(p)$) by $\tilde{T}_{\mathrm{pr}}(p)=\tilde{T}(p)-p^{2}|L\rangle\langle L|$. Since $P(\gamma),\tilde{T}(p)$, and $\tilde{T}_{\mathrm{pr}}(p)$ are all Hermitian and positive definite, and the spectrum of $P(\gamma)\tilde{T}(p)$ is the same as that of $P(\gamma)\tilde{T}_{\mathrm{pr}}(p)$ up to corrections $\sim\mathcal{O}(e^{-\gamma L})$, the spectrum of $P(\gamma)\tilde{T}(p)$ is bounded from above by the product of the largest eigenvalues of $P(\gamma)$ and $\tilde{T}_{\mathrm{pr}}(p)$, which is $e^{-\gamma}4p(1-p)$. We observe numerical evidence that the leading eigenvalue of $P(\gamma)T(p)$ indeed approaches $4p(1-p)$ as $L\to\infty,\gamma\to 0$ (see Fig. 2).
@@
Furthermore, it is straightforward to see that the eigenvector $\overline{\rho_{R,0}}(x)$ of $P(\gamma)T(p)$ corresponding to leading eigenvalue $\lambda_{0}$ is peaked at size $-\frac{\log(\lambda_{0})}{\gamma}$. The eigenstate satisfies
$\begin{split}\lambda_{0}\overline{\rho_{R,k}}(x)&=e^{-\gamma x}\left(2p(1-p)\overline{\rho_{R,0}}(x)+p^{2}\overline{\rho_{R,0}}(x-1)\right)\\ &+e^{-\gamma x}(1-p)^{2}\overline{\rho_{R,0}}(x+1)\end{split}$ (9)
@@
Suppose that $\overline{\rho_{R,0}}(x)$ is a smooth function of $x$ (we will show that this is true in an approximation of $P(\gamma)T(p)$ later in this section; there, $\overline{\rho_{R,0}}$ has a Gaussian-like shape). Then when $\overline{\rho_{R,0}}(x)$ peaks, it flattens out, so $\overline{\rho_{R,0}}(x^{*})\sim\overline{\rho_{R,0}}(x^{*}-1)\sim\overline{\rho_{R,0}}(x^{*}+1)$. If we in addition take $\gamma\to 0$, (9) gives $\lambda_{0}=e^{-\gamma x^{*}}$, so $x^{*}=-\frac{\log{\lambda_{0}}}{\gamma}$. This result is consistent with the argument in Ref. Mori, 2024. Assuming that the leading eigenvalue of $P(\gamma)T(p)$ approaches $4p(1-p)$ in the limit $L\to\infty,\gamma\to 0$, we find that indeed operators get spread to a size $\sim-\frac{\log(4p(1-p))}{\gamma}$. Because the operator size distribution get frozen, and operators do not grow to size $\mathcal{O}(L)$, our argument from the previous section applies to this case as well. Even for Floquet systems, where $T$ may be different from $T(p)$ in (4), we expect the same overall picture to hold: dissipation leads to operators growing ballistically to size $x_{*}=O(1/\gamma)$ and becoming stuck there subsequently, with their norm decaying $e^{-\gamma x_{*}t}$ ever after. It follows from a similar argument to that near Eq. (8) that the dissipative Floquet operator $\mathbb{P}_{\gamma}\mathbb{U}_{F}$ has a leading eigenvalue with modulus $e^{-\bar{g}}$.